Status Report

NASA Space Station On-Orbit Status 26 Nov 2003

By SpaceRef Editor
November 27, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 26 Nov 2003
iss

All ISS systems continue to function nominally, except as noted previously or below. Day 40 in space for Expedition 8 (38 days on ISS).
 

CDR/SO Michael Foale continued his size alteration activity on the LEMS (lower extremity monitoring suit) pants for the FOOT (foot/ground reaction forces during space flight) experiment, using the onboard sewing kit.  The task was started on 11/21 but not completed at that time.  If time permitted, Mike was given the Go to start on the second pair.  [Worn during the actual day-long experiment, the LEMS pants are black Lycra biking tights with 20 electrodes, complemented by shoes fitted with insoles that measure impact forces on the bottom of the feet for the 12-hr session.  Foale will go through a typical on-orbit day while reaction forces against the ISS structure are recorded passively to determine how much stress his legs and feet endure.  This provides better understanding of the bone loss and loss of muscle mass experienced by astronauts in zero-G (on Mir, for example, cosmonauts lost as much bone mass in a month as post-menopausal women do in a year).  The experiment is led by the biomedical engineering department at the Cleveland Clinic Foundation in Cleveland, Ohio.]

Afterwards, Mike Foale conducted a procedure review for the CBOSS-FDI (Cellular Biotechnology Operations Support System/Fluid Dynamics Investigation) experiment, followed by the second of a two-part task determining optimal camera setups for imaging the FDI TCMs (tissue culture modules).  [For today’s part, Mike first made a paper tube for directing light from the flash to the camera in order to contain the amount of flash-produced light reaching the TCM.  Three different scenarios were then imaged: an empty TCM holder, a clear TCM, and a blue TCM (but not all steps were completed today).  Ground specialists were “extremely pleased” with the pictures sent down from the first camera optimization session on 11/20.  They showed very little reflection, if any, and for the first time imaged light transmitted through the sample bag.  As a result, evidence of the effects of microgravity in the fluid could be seen for the first time.]

FE Alexander Kaleri worked in the DC-1 docking compartment, finishing and closing out its restitution to its initial state prior to the Orlan-suited demo last week (11/18).

Kaleri completed the periodic air sampling in the Russian segment (RS), first using the standard Russian AK-1M sampler equipment in the Service Module (SM) and FGB.  Later, testing for CO (carbon monoxide) levels, he collected SM atmospheric samples with the IPD Draeger tubes sampler.  An additional assignment for the FE was the search for and audit of IPD cartridge belts, for updating number and location of the Draeger kits in the inventory management system (IMS) database.

Special instructions were uplinked by TsUP/Moscow for the crew to inspect panels in the area of the 11/24 spill from the leaking EDV-U container with pretreated urine (toxicity level 2).   [If traces were found on the decorative SM panel liner, those liner sections were to be removed.  Ammonia (NH3) concentration in the EDV-U location was to be checked with IPD-1 Draeger tubes.  A second damp cleaning of the spill locations was to be performed, with particular attention on possible contamination of the interior behind panels and hard-to-reach areas.  For the cleaning, Mike and Sasha were told to use wipes saturated with water and dry wipes, and during the cleanup, they were to wear individual protection gear (gloves and goggles).]

Sasha Kaleri completed the daily routine maintenance of the SM’s SOZh life support system, including ASU toilet facilities, as well as the preparation of the “delta” file for updating the IMS, while Mike performed the regular routine status checkup of running Expedition 8 payloads.

The CDR also completed his fourth weekly filling out of the Food Frequency Questionnaire (FFQ), which keeps a log of his nutritional intake over time on the MEC (medical equipment computer).

At 7:40am EST, Foale and Kaleri downlinked a 20-min. TV message with greetings and words of gratitude for Thanksgiving Day, via Ku-band.  It was taped on the ground for replay on NASA TV during the holiday.

TVIS engineers at MCC-H have analyzed the dynamic loads data from the treadmill tests on 11/24 and are working on finalizing a decision on further use of the TVIS for exercising (currently still on hold).  Data were obtained both by the IWIS (interim wireless instrumentation system) and the SAMS/MAMS equipment (space acceleration measurement system/microgravity acceleration measurement system).  [After reviewing the loads data from the walking/running tests without roll-stabilizing gyro and no corner stabilizers, main concern has now shifted from the impacts on the station structure to the long-term status of TVIS hardware.  Additional questions to the crew were uplinked to clarify situational understanding on the ground.  The handrail fit check last weekend showed the TVIS handrail contacting the SM dining table at start-up and slow walking pace, with contact amplitude and frequency decreasing as walking pace increased.  The crew was asked to find some sort of foam material to place between the TVIS and table handrails as a bumper under the SPD (subject positioning device) hook.]

At 10:56am, the ISS performed its attitude transition maneuver from sun-pointing XPOP (x-axis perpendicular to orbit plane) to earth-oriented LVLH (local vertical/local horizontal).  Planning of the maneuver has emphasized conservatism to minimize the gimbal rates on the control moment gyros (CMGs).  [Recap:  On 11/8, CMG-3 exhibited a previously unobserved vibration.  Further analysis revealed other abnormal vibration responses on CMG-3, all associated with CMG desaturations (by Russian thruster firings).  While the direct cause of the vibrations is not understood at this time, it is clear that limiting the gimbal rates associated with desats is less stressful on the CMG bearings and may aid in mitigating the vibration-causing mechanism.]

Early this morning the CDR reported an anomalous “metallic” noise, apparently coming from the outside of the SM.  At about the same time, the ground noted a slight disturbance on the USOS CMGs.  The episode is under investigation; internal pressure monitoring has up to now indicated no cabin air depressurization, nor have any other instruments shown anything amiss.  Later today (~3:00pm EST), the SSRMS (Space Station Remote Manipulator System) will be used in two single-joint maneuvers to do a visual inspection with the elbow tip camera of the general SM location where the noise seemed to originate.

 
Today’s CEO (crew earth observation) targets, limited in XPOP attitude (until 11:00am EST today) by flight rule constraints on the use of the Lab nadir/science window, werePatagonian Glaciers (this was probably the best pass of the day over this target in terms of illumination and weather. There were still some clouds, but the crew was to use the long lenses for detail of the smaller glaciers on the western and eastern flanks of the large Southern Patagonian Ice Field),Rio de Janeiro, Brazil (the crew had a nadir pass with only a few clouds expected over South America’s largest city),Lower Amazon River Basin(the crew had a nadir pass over this target with only fair weather cumulus expected.  Careful mapping of the islands in this large, dynamic estuary will help detect changes in their size and position),Bahamas (a cold front was approaching from the NW on this pass, but the crew still may have been able to get some good views of this picturesque archipelago. We are looking for long-lens low-oblique and near-nadir views of details of the coral reef structures of these islands),Industrialized Southeastern Africa (with high pressure settling over southern Africa, aerosol buildup in the lower atmosphere was likely.  As ISS tracked NE-ward over the interior, the crew was to look right of track in high oblique views for evidence of smog layers over the industrialized Vaal and Orange River valleys), andCongo-Zimbabwe Biomass Burning (as ISS continued NE-ward from southern Africa this pass, the crew was asked to point the camera either side of track for both active fires and burn scars, especially as ISS approached Lake Nyasa).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 19A.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.7; ppO2 (mmHg) — 152.5; ppCO2 (mmHg) — 2.5.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 741.98; temperature (deg C) — 24.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 743.81; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 743.91; temperature (deg C) — 27.0; shell heater temp (deg C) — 23.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.6
  • PMA-2:  Shell heater temp (deg C) — 15.1

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in biased Autotrack mode (suntracking). 
  • SM batteries:  Battery #8, formerly known as #7, is still in slot #8 for troubleshooting; all other batteries (7) are in “Partial Charge” mode (batteries #1 and #3 are degraded).  
  • FGB batteries:  Battery #5 is off; battery #3 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational (all lanes reintegrated 11/5).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available:3704 kg (8166 lb) as of 11/13 [SM(755) + FGB(2597) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH  XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:09am EST [= epoch]):

  • Mean altitude — 374.0 km
  • Apogee — 378.1 km
  • Perigee — 369.9 km
  • Period — 92.0 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006062
  • Orbits per 24-hr. day — 15.65
  • Mean altitude loss in last 24 hours — 185 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28637
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.